Fermi National Laboratory
 

Accelerator Update for February 23, 2009 - February 25, 2009

 

Monday Feb 23

The day shift began with Operations monitoring the Tevatron (TeV) store 6825 (113.45E30) and the Recycler stash (184.2E10), with the Antiproton Source stacking (36.1E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

A MI expert began a MI-8 multiwire study at 10:12 AM.

Operations turned off the Proton Source beam and began preparing Linac and Booster at 11:30 AM for an access at noon.

Operations terminated store 6825 at 12:05 PM.

Linac came out of access at 12:16 PM.

Operations put an energy conservation time line into the Time Line Generator at 12:47 PM.

Operations prepared the TeV, CDF, and D0 for access at 12:54 PM.

CDF came out of access at 1:17 PM.

D0 came out of access at 2:11 PM.

An instrumentation expert reported at 2:12 PM that he was rebooting the MI flying wire system.

Operations reported a problem with the MI beam stability around 3 PM.

Operations reported at 3:15 PM that the TeV access had ended.

Operations established beam to MiniBooNE at 3:35 PM.

At 3:59 PM, Operations reported that the MI beam was stable. A damper problem was causing bad beam positions.

Operations reported at 4:12 PM that the Recycler had stashed antiprotons four times during the day shift.

Operations put the TeV through a dry squeeze at 4:41 PM.

Operations reported at 5:32 PM that they were putting the TeV through the first of three wet squeeze over the next hour.

Operations reported at 6:49 PM that the TeV orbits were good.

Operations put the TeV into shot setup at 6:56 PM.

At 6:57 PM, an Accumulator PLC (A:PLC03) tripped off and caused Pbar Debuncher TWT trips.

At 7:01 PM, Operations reported that every time the TeV took protons from MI, MI beam positions went bad. A TeV expert investigated the energy mismatch.

At 7:33 PM, Operations reported that the MI aborted after every NuMI event. Operations halted the NuMI beam during shot setup.

Operations established store 6829 at 8:30 PM with an initial luminosity of 277.11E30.

Operations reported at 11:51 PM that the Recycler had stashed seven times during the evening shift.

Tuesday Feb 24

The midnight shift began with Operations monitoring store 6829 (160.53E30) and the stash (103.7E10), with Pbar stacking (29.1E10), and with NuMI and MiniBooNE taking beam.

At 1:33 AM, Operations reported that the H- Source output had been dropping. ( Preaccelerator experts switched to the H- Source during the shutdown on Monday.)


Operators reported at 3 AM that they had adjusted the H- Source and improved its output. The plot below showed the magnet current, extractor current, and the Linac output after the adjustment.


At 4:24 AM, LRF3 tripped off. It reset.

At 5:27 AM, two MI kickers (I:KPS4A and B) tripped off. They reset.

Operations reported at 5:48 AM operators had tuned up the 750KeV line.

Operations completed a tune of the 400MeV line at 6:02 AM.

At 7:54 AM, the P1 transfer line tripped off due to a MI power supply (I:V714) drifting low. Everything reset.

At 8:01 AM, the TeV quenched at sector C4. Store 6829 was lost. Operations reported that two TeV quadrupoles (T:QC42 and C:D0Q5) tripped off, but later reported that D0Q2 was the probable cause.

Operations reported at 8:31 AM that the Recycler had stashed eight times during the midnight shift.

Operations prepared TeV sectors A-E for access at 11:02 AM. TeV experts will hunt for a possible ground fault.

A TeV expert reported at 11:51 AM that he had successfully hipotted the D0Q2 magnet, which eliminated the magnet from the ground fault hunt. They began checking the bus cage area.

A Linac expert began a 65-minute calibration of a Linac RF station (LRF3) at 12:12 PM.

At 12:55 PM, the P1 transfer line beam permit tripped off due to I:V714 drifting low. Everything reset.

TeV experts reported at 2:32 PM that the D0Q2 ground fault went away.

MI experts completed their MI-8 wire studies at 2:52 PM.

The TeV came out of access around 3:34 PM. Operations began the TeV turn on.

Operations reported at 4 PM that the Recycler had stashed four times during the day shift.

Operations put the TeV through a dry squeeze at 5:01 PM.

Operations put the TeV through a wet squeeze around 5:56 PM.

Operations put the TeV into shot setup at 7:40 PM.

Operations established store 6832 at 9:07 PM with an initial luminosity of 338.55E30.

Operations reported at 11:59 PM that the Recycler had stashed five times during the evening shift.

Wednesday Feb 25

The midnight shift began with Operations monitoring store 6832 (198.36E30) and the stash (179.2E10), with Pbar stacking (29.8E10), and with NuMI and MiniBooNE taking beam.

At 5:19 AM, Linac Klystron RF station (KRF3) tripped off. It reset.

At 5:27 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 and 8 to also trip. Everything reset.

At 6:20 AM, a MI power supply (I:V714) tripped off. It reset.

At 7:46 AM, KRF5 tripped off. It reset.

Operations reported at 7:58 AM that the Recycler had stashed nine times during the evening shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/04/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory